### v1

# PositionIC: Unified Position and Identity Consistency for Image Customization

Abstract
Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained entity-level spatial control remains elusive, hindering the broader real-world application. This limitation is mainly attributed to the absence of scalable datasets that bind identity with precise positional cues. To this end, we introduce PositionIC, a unified framework that enforces position and identity consistency for multi-subject customization. We construct a scalable synthesis pipeline that employs a bidirectional generation paradigm to eliminate subject drift and maintain semantic coherence. On top of these data, we design a lightweight positional modulation layer that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity. Extensive experiments demonstrate that our approach can achieve precise spatial control while maintaining high consistency in image customization task. PositionIC paves the way for controllable, high-fidelity image customization in open-world, multi-entity scenarios and will be released to foster further research.
Figure 1: Rusults from PositionIC across various controllable image customization tasks.
Introduction
Recently, diffusion-based models have achieved remarkable advances, especially in text-to-image (T2I) generation (Podell et al. 2023; Cai et al. 2025; Rombach et al. 2022; Li et al. 2024, ; Labs 2024). These models can produce images with superior visual quality and high semantic fidelity given short or descriptive prompts.
Among the tasks of image generation, image customization generation has attracted considerable attention in recent years. It requires models to generate images that not only adhere to textual descriptions but also preserve the identity features of reference objects. Although recent methods (Wu et al. 2025; Mou et al. 2025; Tan et al. 2024; Wang et al. 2025) have achieved notable improvements in image fidelity, they still lack customized control over the position, size, and scale of generated subjects and suffer from the limitation of insufficient high-quality customized paired data. While other methods (Wang et al. 2024c; Li et al. 2023) permit the placement of specific objects at specific positions, their position controllability and consistency are still insufficient and underexplored. This limitation hinders the effectiveness in applications where precise spatial arrangement and accurate visual quality are demanded, such as e-commerce product display, storybook illustration, and interior arrangement. As a result, there remains pressing challenges for approaches that can simultaneously ensure high image fidelity and offer flexible, fine-grained control over image attributes.
In this paper, we propose an automatic data curation pipeline and PositionIC, a novel method for controllable image customization to tackle these two challenges, i.e., (1) lack of high-quality customized paired data with position signals; and (2) controllable subject-driven customization guided by the layout. Specifically, regarding the data challenge, we introduce a Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS) pipeline. We utilize a hierarchical training strategy to build a bidirectional generation paradigm from single-object to multi-object and vice versa, which could lift resolution constraints. We adopt Multi-modal Large Language Models (MLLMs) on data filtering to improve data quality. However, MLLMs are prone to visual hallucinations (Han et al. 2024; Pi et al. 2024), which can lead to inaccuracies when comparing visual information. Instead of directly relying on MLLMs to assess the consistency between two images, we carefully design a data filter that leverages expert models to convert visual information into textual descriptions. These descriptions are then used to assist MLLMs in performing multi-dimensional consistency evaluation. For controllable image customization, we propose PositionIC, a lightweight position-controllable framework that decouples layout from subject. It unlocks the fine-grained position control capability of Diffusion Transformers (DiT) through instance- level attention. Additionally, PositionIC does not introduce any additional parameters thus no additional computation is required. By restricting reference image features to focus only on specific regions, PositionIC can generate customized subjects at desired locations, which greatly expands the applications of controllable image customization as shown in Figure 1. Further experiments demonstrate the outstanding universality and remarkable capability of our method.
Our contributions are summarized as follows:
  • We design an automatic framework for data synthesis to obtain high-fidelity paired data, addressing the lack of precise subject consistency and positional control signals in existing public datasets.
  • We propose PositionIC, a lightweight position-controllable framework that decouples layout from subject. It achieves precise placement of multiple subjects through regional attention. At the same time, it explicitly enhances the subject fidelity of Diffusion transformers.
  • Extensive experimental results demonstrate that our method not only achieves the state-of-the-art performances on image customization, but also exhibits the highest precision control capability.
Figure 2: Bidirectional Multi-dimensional Perception Data Synthesis framework. (a) We use Subject200K to train a weak model. (b) Forward generation of multi-subject data pairs. (c) Reverse generation of multi-subject data pairs. (d) We utilize MLLMs to filter out our data pairs.
Figure 3: The overall framework of PositionIC. (a) Reference images and prompts are encoded and concatenated with the latent embeddings $z_{t}$, then the whole token sequence is passed to the double stream blocks of DiT. (b) The mechanism of attention horizon. Each reference image $z_{r}^{i}$ is only visible for the specific area of latent noise $z_{t}$ in the attention map.
Related Work
Subject-driven Generation
In addition to using text prompts for conditional image generation, current diffusion models (Wei et al. 2023; Zhang et al. 2024; Huang et al. 2024; Xiao et al. 2025; Huang et al. 2024) support reference image input to achieve preservation of subject identity. Dreambooth (Ruiz et al. 2023) and LoRa (Hu et al. 2022) control the generation of diffusion models through fine-tuning on the specific subject. IP adapter (Ye et al. 2023) takes the reference images as feature inputs and achieves image-to-image generation by injecting the feature into the cross-attention module of the source model. Recently, DiT-based subject-driven models (Wu et al. 2025; Tan et al. 2024, 2025; Mou et al. 2025; Labs et al. 2025) further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation process.
Position Controllable Generation
Some works (Wang et al. 2024b; Chen et al. 2024; Wang et al. 2024a; Jiménez 2023; Bar-Tal et al. 2023) have attempted to generate with precise layout control but few of them preserving the visual consistency where adhering layout instruction. Gligen (Li et al. 2023) encodes Fourier embedding as grounding token to inject position information. MS-diffusion (Wang et al. 2024c) utilizes grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinders further application.
Method
Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
Customized Data Paired Synthesis
We divide the automated data generation process into three stages. (1) Inspired by UNO, we first train a weak model for image customization tasks using Subject200K (Tan et al. 2024) dataset. The generated results are then segmented and passed through a Flux-Outpainting model with random placement to inject spatial control information. The paired data obtained in this stage is filtered by a filter to ensure high fidelity. These data are then used to train the PositionIC-Single model. (2) In the second stage, we perform forward generation of multi-image data pairs. We independently process Subject200K samples via PositionIC-Single, then randomly pair and position the output as input to the Flux-Outpainting model, thereby obtaining multi-subject paired data. (3) To enhance data diversity and improve generalization, we reverse the above process in the third stage. We first use Multi-modal Large Language Models (MLLMs) to generate text descriptions containing multiple subjects, then employ the Flux model to create high-resolution images. Objects are detected and cropped from these results, individually processed by the PositionIC-Single, resulting in high-resolution multi-subject paired data.
Through bidirectional data synthesis, we construct PIC-400K dataset with noise interference.
Figure 4: Qualitative comparison of single-subject generation with different methods on DreamBench.
Figure 5: Qualitative comparison of multi-subject generation with different methods on DreamBench. We adopt a fixed bounding box (e.g., bottom left and bottom right) for generation.
Figure 6: Inaccuracy of directly using CLIP-I and DINO. The revised score is less affected by the size of the subject which can reflect the subject consistency more authentically.
Multi-dimensional Perception Data Filter
Previous studies have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering after data generation. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color, and details. Lastly, we employ Describe Anything Model (DAM) (Lian et al. 2025), a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
Then we calculate the ranking separately and average them for each image pair to obtain final ranks:
$rank=avg(r(s_{v}),r(s_{vlm}),r(s_{ds})),$ (1)
where $r(\cdot)$ denotes rank. With a lower rank indicating a higher subject similarity.
We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs. The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs.
Regional Attention Horizon
To manage subject consistency generation position simultaneously (e.g., generating the subject in a specific location in the image), we introduce a light-weight approach named attention accumulating, which can unlock DiT’s ability of spatial control without extra training and inference cost. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text attention. If extending to single or multi-subject custom generation via concatenating method like (Wu et al. 2025; Tan et al. 2024; Labs et al. 2025; Zhang et al. 2025), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulating, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon. Previous works (Zhang et al. 2025; Chen et al. 2024) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
As shown in Figure 3, we first encode the reference images $I_{r}^{i}$ to the latent space via VAE $\epsilon(\cdot)$. Encoded results $z_{ref}$ are then concatenated with noise latents $z_{t}$ and text embedding $z_{p}$. This process can be formulated as:
$z_{ref}=[\epsilon(I_{r}^{1}),\epsilon(I_{r}^{2}),...,\epsilon(I_{r}^{N})],$
$z=Concatenate(z_{p},z_{t},z_{ref}),$ (2)
where $N$ is the number of reference images and $z$ denotes the input tokens to the DiT model.
Since the global text prompt contains information about the entire image, generating a reference object in a specific region requires focused and singular attention. They only need to pay attention to the corresponding region while ignoring other irrelevant tokens. Therefore, a limited attention horizon is required. We use a binary attention mask to shield the regions that each token should not directly focus on. As shown in Figure 3, each reference image $z_{ref}$ has a restricted attention horizon, which blocks the attention between itself and other reference images. Moreover, only the area within the bounding box in the noise is visible for the specific reference image. The mechanism of attention horizon mask $M$ can be formulated as:
$M(z_{ref}^{i},z_{ref}^{j})=0,i\neq j,$
$M(z_{ref}^{i},z_{t}^{n})=0,z_{t}^{n}\notin BOX_{i},$
$M(other)=1,$ (3)
where $BOX_{i}$ is the bounding box of $i^{th}$ reference subject in the noise and $z_{t}^{n}$ is the $n^{th}$ patch of noise. Thus, the computation of attention is derived as:
$Attention=Softmax(\frac{QK^{T}}{\sqrt{d}}+\log{M})\cdot V$ (4)
PositionIC-Bench
Most existing customized image evaluation (e.g., DreamBench) lacks explicit spatial position annotations. Thus, there is no universal data benchmark for evaluating subject-driven methods with position control. To address this gap, we propose PositionIC-Bench, a benchmark to evaluate subject consistency and position accuracy simultaneously.
We manually select 252 single-subject samples and 296 multi-subject samples in the benchmark, where the object bounding boxes conform to standard proportions and include challenging positional relationships.
To evaluate our model’s performance on the benchmark, We leverage multi-modal large language models (MLLMs) to detect objects and assess their content, computing the IoU and object category alignment between the generated and reference subjects.
Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench. MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
Experiments
Implementations
Training Detail
Following UNO, we first initialize the model using FLUX.1 dev and apply UnoPE to extend the position embedding of the reference images to the non-overlapping area diagonally. We train a LoRA at the rank of 512 on 8 NVIDIA A100 GPUs and set the total batch size of 128. The learning rate is set to ${10}^{-5}$ with cosine warm up. In the first stage, we train the single-subject model on 44k single-subject pairs for 10k steps. We then continue our training on 54k multi-subject pairs for 8k steps, which extends the multi-subject generation capability to the model obtained in the first stage.
[…]
Dreambooth 0.776 0.215 0.679
BLIP-Diffusion 0.787 0.234 0.742
ELITE 0.803 0.235 0.723
SSR-Encoder 0.797 0.206 0.725
RealCustom 0.783 0.242 0.765
MS-Diffusion 0.808 0.242 0.791
OminiControl 0.805 0.268 0.735
OmniGen 0.791 0.267 0.751
UNO 0.840 0.253 0.814
DreamO 0.835 0.258 0.802
PositionIC$({Ours})$ 0.846 0.269 0.828
Table 1: Evaluation on DreamBench for single-subject driven generation. The bold value is the best and the underlined value is the second best.
[…]
BLIP-Diffusion 0.703 0.212 0.541
MIP-Adapter 0.752 0.254 0.657
MS-Diffusion 0.772 0.261 0.683
OmniGen 0.749 0.291 0.668
UNO 0.781 0.279 0.707
DreamO 0.779 0.273 0.698
PositionIC$({Ours})$ 0.819 0.279 0.771
Table 2: Evaluation on DreamBench for multi-subject driven generation.
Evaluation Metrics
For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench (Ruiz et al. 2023). For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench We use Vision-R1 (Huang et al. 2025) to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
Method Single-Subject Multi-Subject
[…]
Gligen (Li et al. 2023) 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
RegionalFlux (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
MS-Diffusion 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
Instance-Diffusion 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
PositionIC$({Ours})$ 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
Table 3: Quantitative results of controllable spacial generation on PositionIC-Bench.
Qualitative Result
We visualize the comparison results with the current state-of-the-art methods at Figure 4. Overall, our method surpasses all current methods in terms of visual effects. It can be seen that our method can still effectively follow the prompt while maintaining subject consistency, demonstrating higher text fidelity. Other methods either fail to follow complex instructions or cannot maintain consistency. (e.g., UNO and DreamO cannot reproduce the dog face, while SSR-Encoder fails to add the Santa hat.) The results of the third row also reveal that PositionIC has a great capability on patterns and text. Furthermore, PositionIC consistently produces images with a higher degree of naturalness and visual plausibility.
Figure 5 shows the comparison of multi-subject generation. To control variables, the images generated by our method adopt a fixed bounding box (e.g., bottom left and bottom right). It can be seen that in the more difficult multi-subject scenario, PositionIC can still maintain high subject similarity and follow the given text prompt, whereas results from most other methods fail to preserve consistency for each subject and even ignore certain subjects.
We further evaluate the positional control capability of existing methods on PositionIC-Bench via randomly generating the bounding boxes. As shown in Figure 7, PositionIC accurately generates subjects that fully occupy the bounding box without damaging their features. More importantly, our flux-based approach significantly outperforms others in terms of image aesthetic quality and the logical coherence of multi-subject compositions.
Figure 8: Ablation study of data filter. We train our model on PIC-98K, PIC-400K and Subject200K respectively.
Quantitative Evaluations
Subject-driven Analyses
Specifically, we discover that different object sizes can lead to inaccurate scores. As shown in Figure 6, . To avoid the sensitivity, we crop the subject from original image as source images for evaluation.
We compare our proposed PositionIC with several leading methods on DreamBench for both single-subject and multi-subject. As presented in Table 1 and Table 2, PositionIC achieves the highest scores on both CLIP-I and DINO of 0.846 and 0.828 on single-subject, 0.819 and 0.771 on multi-subject, respectively. PositionIC also has competitive CLIP-T compared to existing methods. The evaluation results indicate that PositionIC has remarkable performance on subject consistency and text fidelity.
Controllable Spacial Generation
Table 3 presents the results of spatial control evaluation. PositionIC achieves superior performance in both single-subject and multi-subject position control. For single-subject position control, PositionIC has the highest IoU of 0.828 across all methods and has competitive $AP$ scores compared with Gligen. For multi-subject evaluation, PositionIC achieves the highest scores in both AP and IoU scores, demonstrating a significant advantage over existing methods.
Ablation Study
In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 8 and Table 4.
Impact of data filter. Results in Figure 8 illustrate the advanced fidelity of BMPDS. We directly trained with PIC-98K, PIC-400K and the Subject200K dataset without injecting position control information. PIC-400K achieves significantly higher CLIP-I scores than Subject200K, and the filtered data PIC-98K achieves the highest scores overall, which demonstrates the remarkable efficiency of our data synthesis and filtering pipeline.
[…]
w/o horizon mask 0.784 0.269 0.686
w/ horizon mask 0.846 0.269 0.828
Table 4: Ablation study of horizon mask. Our model perform better subject fidelity on Dreambench after restrict the attention horizon of reference images.
Horizon mask. As shown in Table 4, CLIP-I and DINO scores significantly drop when training without horizon mask. We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistentcy.
Conclusion
In this work, we present PositionIC, an innovative framework capable of customizing multiple subjects with precise position control. PositionIC decouples layout signal from subject feature without introducing additional parameters and training cost. Additionally, we carefully design an automatic data curation framework to obtain high-fidelity paired data. We adopt bidirectional generation and present a multidimensional perception filter to improve object consistency in acquired data. Extensive experiments demonstrate that PositionIC performs high-quality generation in both single-subject and multi-subject consistency, as well as in controllable subject positioning. We hope our work can advance the development of controllable image customization.

### v2

@@
 2024a; Jiménez 2023; Bar-Tal et al.
-2023) have attempted to generate with precise layout control but few of them preserving the visual consistency where adhering layout instruction.
+2023; Shi et al.
+2025) have attempted to generate with precise layout control.
 Gligen (Li et al.
-2023) encodes Fourier embedding as grounding token to inject position information.
+2023) encodes Fourier embedding as grounding tokens to inject position information.
 MS-diffusion (Wang et al.
-2024c) utilizes grounding resampler correlating visual information with specific entities and spatial constraints.
-However, they still suffer from inconsistencies in vision and position, which hinders further application.
+2024c) utilizes a grounding resampler correlating visual information with specific entities and spatial constraints.
+However, they still suffer from inconsistencies in vision and position, which hinder further application.
 Method
@@
 Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench.
+RPF denotes Regional Prompting Flux.
 MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
@@
 DreamO 0.835 0.258 0.802
-PositionIC$({Ours})$ 0.846 0.269 0.828
+PositionIC$({Ours})$ 0.846 0.269 0.823
 Table 1: Evaluation on DreamBench for single-subject driven generation. The bold value is the best and the underlined value is the second best.
@@
 Gligen (Li et al. 2023) 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
-RegionalFlux (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
+RPF (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
 MS-Diffusion 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
@@
 We compare our proposed PositionIC with several leading methods on DreamBench for both single-subject and multi-subject.
-As presented in Table 1 and Table 2, PositionIC achieves the highest scores on both CLIP-I and DINO of 0.846 and 0.828 on single-subject, 0.819 and 0.771 on multi-subject, respectively.
+As presented in Table 1 and Table 2, PositionIC achieves the highest scores on both CLIP-I and DINO of 0.846 and 0.823 on single-subject, 0.819 and 0.771 on multi-subject, respectively.
 PositionIC also has competitive CLIP-T compared to existing methods.
@@
 w/o horizon mask 0.784 0.269 0.686
-w/ horizon mask 0.846 0.269 0.828
+w/ horizon mask 0.846 0.269 0.823
 Table 4: Ablation study of horizon mask. Our model perform better subject fidelity on Dreambench after restrict the attention horizon of reference images.

### v3

@@
 Abstract
-Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained entity-level spatial control remains elusive, hindering the broader real-world application.
+Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained instance-level spatial control remains elusive, hindering the broader real-world application.
 This limitation is mainly attributed to the absence of scalable datasets that bind identity with precise positional cues.
@@
 We construct a scalable synthesis pipeline that employs a bidirectional generation paradigm to eliminate subject drift and maintain semantic coherence.
-On top of these data, we design a lightweight positional modulation layer that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity.
+On top of these data, we design a lightweight positional modulation operation that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity.
 Extensive experiments demonstrate that our approach can achieve precise spatial control while maintaining high consistency in image customization task.
 PositionIC paves the way for controllable, high-fidelity image customization in open-world, multi-entity scenarios and will be released to foster further research.
-Figure 1: Rusults from PositionIC across various controllable image customization tasks.
 Introduction
-Recently, diffusion-based models have achieved remarkable advances, especially in text-to-image (T2I) generation (Podell et al.
-2023; Cai et al.
-2025; Rombach et al.
+Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images that faithfully match short textual prompts (Podell et al.
+2023; Esser et al.
+2024; Rombach et al.
 2022; Li et al.
-2024, ; Labs 2024).
-These models can produce images with superior visual quality and high semantic fidelity given short or descriptive prompts.
-Among the tasks of image generation, image customization generation has attracted considerable attention in recent years.
-It requires models to generate images that not only adhere to textual descriptions but also preserve the identity features of reference objects.
+2024; Labs 2024).
+Within this broader area, subject-driven image customization seeks to generate new scenes that simultaneously (i) conform to user-provided textual descriptions and (ii) preserve the identity of one or more reference objects.
 Although recent methods (Wu et al.
@@
 2024; Wang et al.
-2025) have achieved notable improvements in image fidelity, they still lack customized control over the position, size, and scale of generated subjects and suffer from the limitation of insufficient high-quality customized paired data.
-While other methods (Wang et al.
+2025) markedly improve visual fidelity, they still offer limited control over where or how each subject appears (i.e., position, scale, and relative layout).
+This limitation stems from twofold root causes: scarce data with explicit positional labels and existing global-level attention mechanisms that cannot deliver fine-grained, instance-level guidance.
+Figure 1: Results from PositionIC across various controllable image customization tasks.
+Approaches that do permit object placement (Wang et al.
 2024c; Li et al.
-2023) permit the placement of specific objects at specific positions, their position controllability and consistency are still insufficient and underexplored.
-This limitation hinders the effectiveness in applications where precise spatial arrangement and accurate visual quality are demanded, such as e-commerce product display, storybook illustration, and interior arrangement.
-As a result, there remains pressing challenges for approaches that can simultaneously ensure high image fidelity and offer flexible, fine-grained control over image attributes.
-In this paper, we propose an automatic data curation pipeline and PositionIC, a novel method for controllable image customization to tackle these two challenges, i.e., (1) lack of high-quality customized paired data with position signals; and (2) controllable subject-driven customization guided by the layout.
-Specifically, regarding the data challenge, we introduce a Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS) pipeline.
-We utilize a hierarchical training strategy to build a bidirectional generation paradigm from single-object to multi-object and vice versa, which could lift resolution constraints.
-We adopt Multi-modal Large Language Models (MLLMs) on data filtering to improve data quality.
-However, MLLMs are prone to visual hallucinations (Han et al.
+2023) often compromise either spatial accuracy or identity consistency, restricting their usefulness in practical scenarios such as e-commerce product display, storybook illustration, and interior design.
+Consequently, it remains an open challenge to achieve both high-fidelity rendering and fine-grained, flexible control over subject attributes in a unified framework.
+To this end, we address these issues with two tightly coupled innovations: i) Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS), a scalable pipeline for generating high-quality, position-annotated multi-subject data; ii) PositionIC, a lightweight, layout-aware framework that enables fine-grained, position-controllable, identity-consistent image customization.
+Specifically, to overcome the data bottleneck, we devise an automatic pipeline that expands single-subject collections into scalable, high-quality multi-subject datasets annotated with explicit position masks.
+A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed.
+Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) (Han et al.
 2024; Pi et al.
-2024), which can lead to inaccuracies when comparing visual information.
-Instead of directly relying on MLLMs to assess the consistency between two images, we carefully design a data filter that leverages expert models to convert visual information into textual descriptions.
-These descriptions are then used to assist MLLMs in performing multi-dimensional consistency evaluation.
-For controllable image customization, we propose PositionIC, a lightweight position-controllable framework that decouples layout from subject.
-It unlocks the fine-grained position control capability of Diffusion Transformers (DiT) through instance- level attention.
-Additionally, PositionIC does not introduce any additional parameters thus no additional computation is required.
-By restricting reference image features to focus only on specific regions, PositionIC can generate customized subjects at desired locations, which greatly expands the applications of controllable image customization as shown in Figure 1.
-Further experiments demonstrate the outstanding universality and remarkable capability of our method.
+2024), we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks.
+This two-stage filtering markedly improves data reliability.
+Built upon the BMPDS corpus, PositionIC injects an instance-level attention modulation into Diffusion Transformer (DiT), endowing its fine-grained position control capability.
+By decoupling instance-level spatial embeddings from semantic identity features, our method enables independent, accurate placement of each subject while introducing no extra train-time parameters or inference overhead.
+Restricting reference features to user-specified regions further enhances identity fidelity and spatial precision, unlocking new applications of controllable image customization (Figure 1).
 Our contributions are summarized as follows:
@@
 In addition to using text prompts for conditional image generation, current diffusion models (Wei et al.
-2023; Zhang et al.
-2024; Huang et al.
-2024; Xiao et al.
-2025; Huang et al.
-2024) support reference image input to achieve preservation of subject identity.
+2023; Li, Li, and Hoi 2023; Huang et al.
+2024; Ye et al.
+2023; Xiong et al.
+2025; Xiao et al.
+2023; Bar-Tal et al.
+2023; Feng et al.
+2025) support reference image input to achieve preservation of subject identity.
 Dreambooth (Ruiz et al.
@@
 2022) control the generation of diffusion models through fine-tuning on the specific subject.
-IP adapter (Ye et al.
-2023) takes the reference images as feature inputs and achieves image-to-image generation by injecting the feature into the cross-attention module of the source model.
-Recently, DiT-based subject-driven models (Wu et al.
+Recently, Diffisuion-Transformers-based subject-driven models (Wu et al.
 2025; Tan et al.
 2024, 2025; Mou et al.
+2025; Xiao et al.
 2025; Labs et al.
@@
 Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals.
-However, existing open-source datasets such as Subject200K generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information.
+However, existing open-source datasets such as Subject200K (Tan et al.
+2024) generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information.
 We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations.
@@
 We divide the automated data generation process into three stages.
-(1) Inspired by UNO, we first train a weak model for image customization tasks using Subject200K (Tan et al.
-2024) dataset.
+(1) Inspired by UNO, we first train a weak model for image customization tasks using Subject200K dataset.
 The generated results are then segmented and passed through a Flux-Outpainting model with random placement to inject spatial control information.
@@
 (3) To enhance data diversity and improve generalization, we reverse the above process in the third stage.
-We first use Multi-modal Large Language Models (MLLMs) to generate text descriptions containing multiple subjects, then employ the Flux model to create high-resolution images.
+We first use Large Language Models (LLMs) to generate text descriptions containing multiple subjects, then employ the Flux model to create high-resolution images.
 Objects are detected and cropped from these results, individually processed by the PositionIC-Single, resulting in high-resolution multi-subject paired data.
-Through bidirectional data synthesis, we construct PIC-400K dataset with noise interference.
+Through bidirectional data synthesis, we construct PIC-400K dataset. However, we need extra filtering processes to improve data quality as the dataset is still suffering from noise.
 Figure 4: Qualitative comparison of single-subject generation with different methods on DreamBench.
@@
 Multi-dimensional Perception Data Filter
-Previous studies have shown that MLLMs have limited capability in recognizing fine-grained details in images.
+Previous studies (Cao et al.
+2024; He et al.
+2025) have shown that MLLMs have limited capability in recognizing fine-grained details in images.
 Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering after data generation.
@@
 where $r(\cdot)$ denotes rank. With a lower rank indicating a higher subject similarity.
-We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs. The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs.
+We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs.
+The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs.
+More examples of our PIC-98K are depicted in Appendices.
 Regional Attention Horizon
@@
 2025; Zhang et al.
-2025), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
-To achieve effective attention accumulating, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon.
+2025a), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
+To achieve effective attention accumulation, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon.
 Previous works (Zhang et al.
-2025; Chen et al.
-2024) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
+2025a; Chen et al.
+2024; Zhang et al.
+2025b) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
 As shown in Figure 3, we first encode the reference images $I_{r}^{i}$ to the latent space via VAE $\epsilon(\cdot)$.
@@
 We manually select 252 single-subject samples and 296 multi-subject samples in the benchmark, where the object bounding boxes conform to standard proportions and include challenging positional relationships.
-To evaluate our model’s performance on the benchmark, We leverage multi-modal large language models (MLLMs) to detect objects and assess their content, computing the IoU and object category alignment between the generated and reference subjects.
 Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench.
-RPF denotes Regional Prompting Flux.
+MS-Diff and RPF denote MS-Diffusion and Regional Prompting Flux respectively.
 MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
@@
 MS-Diffusion 0.808 0.242 0.791
+OmniGen 0.791 0.267 0.751
+DreamO 0.835 0.258 0.802
 OminiControl 0.805 0.268 0.735
-OmniGen 0.791 0.267 0.751
 UNO 0.840 0.253 0.814
-DreamO 0.835 0.258 0.802
-PositionIC$({Ours})$ 0.846 0.269 0.823
-Table 1: Evaluation on DreamBench for single-subject driven generation. The bold value is the best and the underlined value is the second best.
+PositionIC(Ours)({Ours}) 0.846 0.269 0.823
+Table 1: Evaluation on DreamBench for single-subject driven generation. The bold value is the highest and the underlined value is the second.
 […]
@@
 MS-Diffusion 0.772 0.261 0.683
+DreamO 0.779 0.273 0.698
+UNO 0.781 0.279 0.707
 OmniGen 0.749 0.291 0.668
-UNO 0.781 0.279 0.707
-DreamO 0.779 0.273 0.698
-PositionIC$({Ours})$ 0.819 0.279 0.771
+PositionIC(Ours)({Ours}) 0.819 0.279 0.771
 Table 2: Evaluation on DreamBench for multi-subject driven generation.
@@
 For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
-For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench We use Vision-R1 (Huang et al.
+For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench.
+We use Vision-R1 (Huang et al.
 2025) to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
@@
 […]
+RPF (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
+MS-Diffusion (Wang et al. 2024c) 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
+Instance-Diffusion (Wang et al. 2024b) 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
 Gligen (Li et al. 2023) 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
-RPF (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
-MS-Diffusion 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
-Instance-Diffusion 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
-PositionIC$({Ours})$ 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
+PositionIC(Ours)({Ours}) 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
 Table 3: Quantitative results of controllable spacial generation on PositionIC-Bench.
 Qualitative Result
-We visualize the comparison results with the current state-of-the-art methods at Figure 4.
+We visualize the comparison results with the current state-of-the-art methods in Figure 4.
 Overall, our method surpasses all current methods in terms of visual effects.
@@
 To control variables, the images generated by our method adopt a fixed bounding box (e.g., bottom left and bottom right).
-It can be seen that in the more difficult multi-subject scenario, PositionIC can still maintain high subject similarity and follow the given text prompt, whereas results from most other methods fail to preserve consistency for each subject and even ignore certain subjects.
+In a more difficult multi-subject scenario, PositionIC can still maintain high subject similarity and follow the given text prompt, whereas results from most other methods fail to preserve consistency for each subject and even ignore certain subjects.
 We further evaluate the positional control capability of existing methods on PositionIC-Bench via randomly generating the bounding boxes.
@@
 More importantly, our flux-based approach significantly outperforms others in terms of image aesthetic quality and the logical coherence of multi-subject compositions.
+Figure 8: User study for subject-driven generation and data filter.
+(a) User evaluation on Dreambench.
+(b) Filtering consistency of our data filter compared with human in different data quality intervals (from good to bad).
+Quantitative Evaluations
 Figure 9: Ablation study of data filter. We train our model on PIC-98K, PIC-400K and Subject200K respectively.
-Quantitative Evaluations
 Subject-driven Analyses
 Specifically, we discover that different object sizes can lead to inaccurate scores.
-As shown in Figure 6, .
-To avoid the sensitivity, we crop the subject from original image as source images for evaluation.
+As shown in Figure 6, to avoid the sensitivity, we crop the subject from original image as source images for evaluation.
 We compare our proposed PositionIC with several leading methods on DreamBench for both single-subject and multi-subject.
@@
 For multi-subject evaluation, PositionIC achieves the highest scores in both AP and IoU scores, demonstrating a significant advantage over existing methods.
+User Study
+We invite evaluators for an extensive user study.
+For subject-driven generation, we randomly selected 500 images from the results on DreamBench for manual evaluation.
+There are six users scored the results from five dimensions, with each ranging from 0 to 5, and the average score was taken.
+The result presented in Figure 8 (a) shows that our method reaches the highest capability to preserve image features and details while maintaining competitive text adherence.
+To evaluate the consistency between human annotators and the data filter in BMPDS, we use percentage agreement metric (denoted as Acc.), comparing the filter’s output against human-generated annotations.
+As shown in Figure 8 (b), our filter has an average consistency of 0.89 with human annotators.
 Ablation Study
-In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 8 and Table 4.
+In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 9, Table 4 and Figure 10.
 Impact of data filter.
@@
 We directly trained with PIC-98K, PIC-400K and the Subject200K dataset without injecting position control information.
-PIC-400K achieves significantly higher CLIP-I scores than Subject200K, and the filtered data PIC-98K achieves the highest scores overall, which demonstrates the remarkable efficiency of our data synthesis and filtering pipeline.
+PIC-400K achieves significantly higher scores than Subject200K, and the filtered data PIC-98K achieves the highest scores overall, which demonstrates the remarkable efficiency of our data synthesis and filtering pipeline.
 […]
@@
 Table 4: Ablation study of horizon mask. Our model perform better subject fidelity on Dreambench after restrict the attention horizon of reference images.
+Figure 10: Qualitative results of ablation study. Horizon mask and our filter pipeline are capable of effective position control and consistent customization.
 Horizon mask.
 As shown in Table 4, CLIP-I and DINO scores significantly drop when training without horizon mask.
-We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistentcy.
+We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistency.
 Conclusion
@@
 We hope our work can advance the development of controllable image customization.
+Appendix A Bidirectional Multi-dimensional Perception Data Synthesis
+Detailed Instruction of GPT-4o
+As shown in Figure 14, the message given to GPT-4o consists of Instruction, Evaluation Metric and Response.
+In the part of Instruction, we have defined the input format and evaluation metric dimensions, and required GPT-4o to select no fewer than three features for scoring based on the content of textual description.
+In the next part of Evaluation Metric, we detail the metric standard and provide GPT-4o examples to evaluate.
+There are 6 levels, ranging from 0 to 5, representing the similarity between two descriptions regarding the same feature.
+After that, we prompt GPT-4o to return a dictionary in JSON format containing the subject types and the scores for each feature.
+If there is no similarity between the two descriptions, the subject is set to ”none”, indicating that the final score is 0.
+Due to the substantial differences in textual descriptions between subjects, it is not feasible to predetermine the feature categories for evaluation.
+Therefore, we allow the LLMs to select at least three features and assign individual scores to each.
+The final score is calculated as the average of all feature scores.
+For descriptions with significant discrepancies, the LLMs is permitted to assign a score of zero to the samples.
+Figure 15 demonstrates the samples of Multi-dimensional Perception Data Filter.
+We have highlighted the correlated features in the description.
+In the first sample, the teddy bear share the same physical characteristics except for their posture, hence earning the highest appearance score and slightly lower posture score.
+In the second sample, the deer is missing antlers, which resulted in the lowest score on the ”antler” feature.
+Details of PIC-98K Dataset
+We propose PIC-400K utilizing our Bidirectional Multi-dimensional Perception Data Synthesis, a automatic and effective high-consistency data synthesis pipeline.
+Samples of the filtered data PIC-98K is shown in the Figure 13.
+BMPDS can synthesize high-fidelity multi-subject images while maintaining high resolution.
+Against previous works, the position of subjects is controllable and it is randomly placed to train the position control capability of PositionIC.
+There are over 9000 subject descriptions in PIC-98K, including multiple categories such as fruits, animals, and transportation vehicles, which basically cover common objects.
+The distribution of subjects is shown in Figure 11, vehicles, furniture, animal, and kitchenware constitute a significant proportion, with most difficult-to-classify subjects categorized as ”other”.
+Figure 12: Showcases of PositionIC-Bench.
+Appendix B PositionIC-Bench
+We manually select 252 single-subject samples and 296 multi-subject samples in the benchmark, where the object bounding boxes conform to standard proportions and include challenging positional relationships.
+We show some samples of PositionIC-Bench in Figure 12.
+Our bench includes various subjects such as furniture, animals, plants, and portraits.
+Not limited to conventional object placement, PositionIC-Bench’s bounding boxes have more complex spatial relationships where objects are placed on different planes.
+At the same time, the bounding boxes of smaller objects is appropriately enlarged to obtain more accurate evaluation scores.
+Figure 13: Showcases of PIC-98K Dataset.
+Figure 14: Prompt template of MLLMs in Multi-dimensional Perception Data Filter.
+Figure 15: Examples of Multi-dimensional Perception Data Filter.

### v4

@@
 Abstract
-Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained instance-level spatial control remains elusive, hindering the broader real-world application.
-This limitation is mainly attributed to the absence of scalable datasets that bind identity with precise positional cues.
-To this end, we introduce PositionIC, a unified framework that enforces position and identity consistency for multi-subject customization.
-We construct a scalable synthesis pipeline that employs a bidirectional generation paradigm to eliminate subject drift and maintain semantic coherence.
-On top of these data, we design a lightweight positional modulation operation that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity.
-Extensive experiments demonstrate that our approach can achieve precise spatial control while maintaining high consistency in image customization task.
-PositionIC paves the way for controllable, high-fidelity image customization in open-world, multi-entity scenarios and will be released to foster further research.
-Introduction
-Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images that faithfully match short textual prompts (Podell et al.
-2023; Esser et al.
-2024; Rombach et al.
-2022; Li et al.
-2024; Labs 2024).
-Within this broader area, subject-driven image customization seeks to generate new scenes that simultaneously (i) conform to user-provided textual descriptions and (ii) preserve the identity of one or more reference objects.
-Although recent methods (Wu et al.
-2025; Mou et al.
-2025; Tan et al.
-2024; Wang et al.
-2025) markedly improve visual fidelity, they still offer limited control over where or how each subject appears (i.e., position, scale, and relative layout).
-This limitation stems from twofold root causes: scarce data with explicit positional labels and existing global-level attention mechanisms that cannot deliver fine-grained, instance-level guidance.
+Recent subject-driven image customization excels in fidelity, yet fine-grained instance-level spatial control remains an elusive challenge, hindering real-world applications.
+This limitation stems from two factors: a scarcity of scalable, position-annotated datasets, and the entanglement of identity and layout by global attention mechanisms.
+To this end, we introduce PositionIC, a unified framework for high-fidelity, spatially controllable multi-subject customization.
+First, we present BMPDS, the first automatic data-synthesis pipeline for position-annotated multi-subject datasets, effectively providing crucial spatial supervision.
+Second, we design a lightweight, layout-aware diffusion framework that integrates a novel visibility-aware attention mechanism.
+This mechanism explicitly models spatial relationships via an NeRF-inspired volumetric weight regulation to effectively decouple instance-level spatial embeddings from semantic identity features, enabling precise, occlusion-aware placement of multiple subjects.
+Extensive experiments demonstrate PositionIC achieves state-of-the-art performance on public benchmarks, setting new records for spatial precision and identity consistency.
+Our work represents a significant step towards truly controllable, high-fidelity image customization in multi-entity scenarios.
+Code and data will be publicly released.
+1 Introduction
+Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images closely aligned with short textual prompts [sdxl, stablediffusion3, ldm, playground, flux].
+Beyond generic generation, subject-driven image customization aims to integrate one or more reference objects into user-specified scenes under the guidance of a text prompt.
+Traditionally, this task emphasizes two core goals: (i) preserving the appearance identity of the reference object(s), and (ii) aligning the synthesis with the user’s textual description, where remarkable improvements in visual fidelity have been achieved by recent methods [uno, dreamo, omini_control, wang2025unicombine].
 Figure 1: Results from PositionIC across various controllable image customization tasks.
-Approaches that do permit object placement (Wang et al.
-2024c; Li et al.
-2023) often compromise either spatial accuracy or identity consistency, restricting their usefulness in practical scenarios such as e-commerce product display, storybook illustration, and interior design.
-Consequently, it remains an open challenge to achieve both high-fidelity rendering and fine-grained, flexible control over subject attributes in a unified framework.
-To this end, we address these issues with two tightly coupled innovations: i) Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS), a scalable pipeline for generating high-quality, position-annotated multi-subject data; ii) PositionIC, a lightweight, layout-aware framework that enables fine-grained, position-controllable, identity-consistent image customization.
+However, through exploring real-world applications like e-commerce product display, storybook illustration, and interior design, we identified an additional critical dimension: fine-grained, flexible spatial control, encompassing aspects such as subject placement, scale, occlusion, and relative layout.
+In other words, while prior works focus on “what” to generate, practical usage demands where and how each subject appears.
+To explicitly address this, we adopt a refined task definition: subject-driven customization with both identity fidelity and instance-level spatial controllability.
+This is a natural extension of the established task, not a fundamentally different objective, but one that better matches application requirements.
+Despite substantial progress in identity-preserving personalization methods, current approaches still struggle with flexible placement and layout control.
+Specifically, two coupled bottlenecks remain: (1) the scarcity of large-scale datasets annotated with explicit positional labels of multiple subjects, hindering the learning of spatial reasoning; and (2) the prevalent use of global-level attention mechanisms in diffusion transformers, which inadvertently entangle semantic identity and spatial layout, thereby failing to provide fine-grained, visibility-aware control.
+While a few recent works [ms-diffusion, gligen] have attempted layout control, they typically exhibit a trade-off between spatial precision and subject identity consistency, thus limiting their practical utility.
+To this end, we introduce two tightly coupled components: (i) a scalable Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS) pipeline that generates high-quality, multi-subject, position-annotated data; and (ii) PositionIC, a lightweight layout-aware diffusion framework incorporating a visibility-aware attention mechanism that enables precise, controllable placement of multiple subjects while maintaining identity fidelity.
 Specifically, to overcome the data bottleneck, we devise an automatic pipeline that expands single-subject collections into scalable, high-quality multi-subject datasets annotated with explicit position masks.
 A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed.
-Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) (Han et al.
-2024; Pi et al.
-2024), we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks.
+Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) [han2024instinctive, bpo], we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks.
 This two-stage filtering markedly improves data reliability.
-Built upon the BMPDS corpus, PositionIC injects an instance-level attention modulation into Diffusion Transformer (DiT), endowing its fine-grained position control capability.
-By decoupling instance-level spatial embeddings from semantic identity features, our method enables independent, accurate placement of each subject while introducing no extra train-time parameters or inference overhead.
+Built upon the BMPDS corpus, PositionIC directly integrates our visibility-aware attention mechanism into the diffusion transformer architecture.
+This mechanism is achieved mainly via an operation inspired by NeRF [nerf] termed Volumetric Weight Regulation, which explicitly constructs attention masks based on the likelihood of visibility derived from volumetric rendering principles.
+Unlike prevalent global-level attention, our visibility-aware attention mechanism effectively decouples instance-level spatial embeddings from semantic identity features.
+This design enables independent and accurate placement of each subject by explicitly modeling occlusion and perspective relationships, and crucially introduces no extra train-time parameters or inference overhead.
 Restricting reference features to user-specified regions further enhances identity fidelity and spatial precision, unlocking new applications of controllable image customization (Figure 1).
 Our contributions are summarized as follows:
-  • We design an automatic framework for data synthesis to obtain high-fidelity paired data, addressing the lack of precise subject consistency and positional control signals in existing public datasets.
-  • We propose PositionIC, a lightweight position-controllable framework that decouples layout from subject.
-It achieves precise placement of multiple subjects through regional attention.
-At the same time, it explicitly enhances the subject fidelity of Diffusion transformers.
-  • Extensive experimental results demonstrate that our method not only achieves the state-of-the-art performances on image customization, but also exhibits the highest precision control capability.
+  • We present BMPDS, the first large-scale automatic data‐synthesis framework for high-fidelity, position-annotated multi‐subject pairs, filling a critical gap in spatial supervision.
+  • We propose PositionIC, a lightweight yet highly effective layout-aware diffusion framework.
+It integrates a novel visibility-aware attention mechanism into the diffusion transformer architecture.
+This design explicitly decouples spatial layout from semantic identity, enabling unprecedentedly accurate multi-subject placement and occlusion-aware rendering without additional training parameters or inference overhead.
+  • Extensive experimental results demonstrate that our approach delivers state-of-the-art performance on subject-driven customization benchmarks, achieving the highest spatial precision and identity consistency to date.
 Figure 2: Bidirectional Multi-dimensional Perception Data Synthesis framework.
@@
 (d) We utilize MLLMs to filter out our data pairs.
-Figure 3: The overall framework of PositionIC.
-(a) Reference images and prompts are encoded and concatenated with the latent embeddings $z_{t}$, then the whole token sequence is passed to the double stream blocks of DiT.
-(b) The mechanism of attention horizon.
-Each reference image $z_{r}^{i}$ is only visible for the specific area of latent noise $z_{t}$ in the attention map.
-Related Work
-Subject-driven Generation
-In addition to using text prompts for conditional image generation, current diffusion models (Wei et al.
-2023; Li, Li, and Hoi 2023; Huang et al.
-2024; Ye et al.
-2023; Xiong et al.
-2025; Xiao et al.
-2023; Bar-Tal et al.
-2023; Feng et al.
-2025) support reference image input to achieve preservation of subject identity.
-Dreambooth (Ruiz et al.
-2023) and LoRa (Hu et al.
-2022) control the generation of diffusion models through fine-tuning on the specific subject.
-Recently, Diffisuion-Transformers-based subject-driven models (Wu et al.
-2025; Tan et al.
-2024, 2025; Mou et al.
-2025; Xiao et al.
-2025; Labs et al.
-2025) further advance subject-driven generation.
-They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation process.
-Position Controllable Generation
-Some works (Wang et al.
-2024b; Chen et al.
-2024; Wang et al.
-2024a; Jiménez 2023; Bar-Tal et al.
-2023; Shi et al.
-2025) have attempted to generate with precise layout control.
-Gligen (Li et al.
-2023) encodes Fourier embedding as grounding tokens to inject position information.
-MS-diffusion (Wang et al.
-2024c) utilizes a grounding resampler correlating visual information with specific entities and spatial constraints.
+Figure 3: The overall framework of PositionIC.(a) Reference images and prompts are encoded and concatenated with the latent embeddings $z_{t}$, then the whole token sequence is passed to DiT.
+Each reference image $z_{ref}^{i}$ is only visible for the specific area of latent noise $z_{t}$ in the attention map.
+(b) The objects’ binary mask and the semantic density $\sigma$ are used by VWR to calculate the weighted mask.
+The dashed boxes in the ray diagram represent the overlapping regions, where the same colors correspond to the weights in the mask.
+2 Related Work
+2.1 Subject-driven Generation
+In addition to using text prompts for conditional image generation, current diffusion models [ELITE, li2023blip, huang2024realcustom, ip-adapter, groundingbooth, fastcomposer, MultiDiffusion, feng2025personalize] support reference image input to achieve preservation of subject identity.
+Dreambooth [dreambooth] and LoRa [lora] control the generation of diffusion models through fine-tuning on the specific subject.
+Recently, Diffisuion-Transformers-based subject-driven models [uno, omini_control, omini_control2, dreamo, xiao2025omnigen, flux-kontext] further advance subject-driven generation.
+They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation.
+2.2 Position Controllable Generation
+Some works [instancediffusion, regional_flux, wang2024compositional, jimenez2023mixture, MultiDiffusion, shi2025layoutcot, LaRender] have attempted to generate with precise layout control.
+Gligen [gligen] encodes Fourier embedding as grounding tokens to inject position information.
+MS-diffusion [ms-diffusion] utilizes a grounding resampler correlating visual information with specific entities and spatial constraints.
 However, they still suffer from inconsistencies in vision and position, which hinder further application.
-Method
-Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
+3 Method
+3.1 Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
 Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals.
-However, existing open-source datasets such as Subject200K (Tan et al.
-2024) generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information.
+However, existing open-source datasets such as Subject200K [omini_control] generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information.
 We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations.
@@
 The overall framework is depicted in Figure 2.
-Customized Data Paired Synthesis
+3.1.1 Customized Data Paired Synthesis
 We divide the automated data generation process into three stages.
@@
 Through bidirectional data synthesis, we construct PIC-400K dataset. However, we need extra filtering processes to improve data quality as the dataset is still suffering from noise.
-Figure 4: Qualitative comparison of single-subject generation with different methods on DreamBench.
-Figure 5: Qualitative comparison of multi-subject generation with different methods on DreamBench. We adopt a fixed bounding box (e.g., bottom left and bottom right) for generation.
-Figure 6: Inaccuracy of directly using CLIP-I and DINO. The revised score is less affected by the size of the subject which can reflect the subject consistency more authentically.
-Multi-dimensional Perception Data Filter
-Previous studies (Cao et al.
-2024; He et al.
-2025) have shown that MLLMs have limited capability in recognizing fine-grained details in images.
-Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering after data generation.
+3.1.2 Multi-dimensional Perception Data Filter
+Previous studies [cao2024mmfuser, he2025analyzing] have shown that MLLMs have limited capability in recognizing fine-grained details in images.
+Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering.
 Specifically, we divide the filtering process into three levels based on granularity.
-Firstly, we segment the subjects from data pairs and utilize CLIP-I (Radford et al.
-2021) and DINO (Caron et al.
-2021) scores $s_{v}$ to filter out images with significantly lower consistency.
-Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color, and details.
-Lastly, we employ Describe Anything Model (DAM) (Lian et al.
-2025), a description expert to obtain detailed textual descriptions for each subject.
+Firstly, we segment the subjects from data pairs and utilize CLIP-I [CLIP] and DINO [DINO] scores $s_{v}$ to filter out images with significantly lower consistency.
+Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color and so on.
+Lastly, we employ Describe Anything Model (DAM) [describe_anything], a description expert to obtain detailed textual descriptions for each subject.
 Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@
 The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs.
-More examples of our PIC-98K are depicted in Appendices.
-Regional Attention Horizon
-To manage subject consistency generation position simultaneously (e.g., generating the subject in a specific location in the image), we introduce a light-weight approach named attention accumulating, which can unlock DiT’s ability of spatial control without extra training and inference cost.
-As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text attention.
-If extending to single or multi-subject custom generation via concatenating method like (Wu et al.
-2025; Tan et al.
-2024; Labs et al.
-2025; Zhang et al.
-2025a), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
-To achieve effective attention accumulation, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon.
-Previous works (Zhang et al.
-2025a; Chen et al.
-2024; Zhang et al.
-2025b) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
+More examples of our PIC-98K are depicted in the appendix.
+3.2 PositionIC
+Figure 4: Qualitative comparison of single-subject generation with different methods on DreamBench.
+Figure 5: Qualitative comparison of multi-subject generation with different methods on DreamBench. We adopt a fixed bounding box (e.g., bottom left and bottom right) for generation.
+When extending text-to-image models to image-to-image tasks, the fidelity of the generated object largely depends on the magnitude of the attention between the corresponding regions in the target image and the reference image.
+We refer to this as attention accumulation.
+As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text cross attention.
+If extending to single or multi-subject custom generation via concatenating method like [uno, omini_control, flux-kontext, eligen], the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
+To achieve effective attention accumulation, we propose Visibility-Aware Attention, which utilizes volume rendering(VR) to explicitly define the area that can be focused on for each reference image.
+Previous works [eligen, regional_flux, zhang2025creatidesign, LaRender] explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
+The overall mechanism is shown in Figure 3.
+3.2.1 Volume Rendering(VR)
+The weighted mask utilizes VR [volume_render] to handle the occlusion relationships between individual objects.
+Similar to NeRF [nerf], the expected color $C(r)$ observed by the camera along the distance from the far bound $t_{f}$ to the near bound $t_{n}$ is:
+$C(r)=\int_{t_{n}}^{t_{f}}T(t)\sigma(r(t))c(r(t),d)dt,$
+$T(t)=exp(-\int_{t_{n}}^{t}\sigma(r(s))ds)$ (2)
+where the camera ray $r(t)=o+td$, $\sigma(x)$ is volume density. We utilize the quadrature rule to obtain a discrete estimate of $C(\mathbf{r})$:
+$\hat{C}(r)=\sum_{i=1}^{N}T_{i}(1-exp(-\sigma_{i}\delta_{i}))c_{i},$
+$T_{i}=exp(-\sum_{j=1}^{i-1}\sigma_{i}\delta_{i})$ (3)
+where $N$ denotes the spatial resolution for the distance range from near to far bounds. $\delta_{i}$ is the distance between adjacent samples.
+3.2.2 Volumetric Weight Regulation
+Similar to the 3D rendering process, controlling the generation position of individual objects on a 2D canvas can be viewed as a composite image captured by an orthographic virtual camera.
+To precisely distinguish the occlusion relationships between objects, we leverage VR to modulate a physically plausible semantic mask.
+Given the foreground mask $M_{i}$ and the sample distance $\delta$, the volumetric weight $\hat{M_{n}}$ can be formulate as:
+$\hat{M}_{n}=exp(-\sum_{j=1}^{n-1}\sigma_{i}\delta)(1-exp(-\sigma_{n}\delta))M_{n}\hat{M}_{n-1}$ (4)
+unlike in  Sec.
+3.2.1, $\hat{M_{n}}$ is defined as the volumetric weight for each distance level, which is utilized to represent the degree of attention the image pays to each region of the reference image during attention calculation.
+Specifically, we utilize a semantic density $\sigma_{i}$ instead of the volume density to represent the semantic interaction between objects.
+Setting of $\sigma_{i}$ can be found in appendix.
+3.2.3 Visibility-Aware Attention
 As shown in Figure 3, we first encode the reference images $I_{r}^{i}$ to the latent space via VAE $\epsilon(\cdot)$.
@@
 They only need to pay attention to the corresponding region while ignoring other irrelevant tokens.
-Therefore, a limited attention horizon is required.
-We use a binary attention mask to shield the regions that each token should not directly focus on.
+Therefore, we use a VWR-based attention mask, which we call visibility-aware attention(VAA), to shield the regions that each token should not directly focus on.
 As shown in Figure 3, each reference image $z_{ref}$ has a restricted attention horizon, which blocks the attention between itself and other reference images.
 Moreover, only the area within the bounding box in the noise is visible for the specific reference image.
-The mechanism of attention horizon mask $M$ can be formulated as:
+The mechanism of VAA mask $M$ can be formulated as:
 $M(z_{ref}^{i},z_{ref}^{j})=0,i\neq j,$
-$M(z_{ref}^{i},z_{t}^{n})=0,z_{t}^{n}\notin BOX_{i},$
+$M(z_{ref}^{i},z_{t}^{n})=\hat{M}_{i},$
 $M(other)=1,$ (6)
-where $BOX_{i}$ is the bounding box of $i^{th}$ reference subject in the noise and $z_{t}^{n}$ is the $n^{th}$ patch of noise. Thus, the computation of attention is derived as:
+where $\hat{M}_{i}$ is volumetric weight of $i^{th}$ reference subject. Thus, the computation of attention is derived as:
 $Attention=Softmax(\frac{QK^{T}}{\sqrt{d}}+\log{M})\cdot V$ (7)
-PositionIC-Bench
+3.3 PositionIC-Bench
 Most existing customized image evaluation (e.g., DreamBench) lacks explicit spatial position annotations.
@@
 MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
-Experiments
-Implementations
-Training Detail
+4 Experiments
+4.1 Implementations
+4.1.1 Training Detail
 Following UNO, we first initialize the model using FLUX.1 dev and apply UnoPE to extend the position embedding of the reference images to the non-overlapping area diagonally.
@@
 Table 2: Evaluation on DreamBench for multi-subject driven generation.
-Evaluation Metrics
-For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench (Ruiz et al.
-2023).
+4.1.2 Evaluation Metrics
+For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench [dreambooth].
 For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
 For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench.
-We use Vision-R1 (Huang et al.
-2025) to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
+We use Vision-R1 [visionr1] to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
 Method Single-Subject Multi-Subject
 […]
-RPF (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
-MS-Diffusion (Wang et al. 2024c) 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
-Instance-Diffusion (Wang et al. 2024b) 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
-Gligen (Li et al. 2023) 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
+RPF [regional_flux] 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
+MS-Diffusion [ms-diffusion] 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
+Instance-Diffusion [instancediffusion] 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
+Gligen [gligen] 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
 PositionIC(Ours)({Ours}) 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
 Table 3: Quantitative results of controllable spacial generation on PositionIC-Bench.
-Qualitative Result
+4.2 Qualitative Result
 We visualize the comparison results with the current state-of-the-art methods in Figure 4.
@@
 Furthermore, PositionIC consistently produces images with a higher degree of naturalness and visual plausibility.
+More examples can be found in Appendix.
 Figure 5 shows the comparison of multi-subject generation.
@@
 More importantly, our flux-based approach significantly outperforms others in terms of image aesthetic quality and the logical coherence of multi-subject compositions.
-Figure 8: User study for subject-driven generation and data filter.
-(a) User evaluation on Dreambench.
-(b) Filtering consistency of our data filter compared with human in different data quality intervals (from good to bad).
-Quantitative Evaluations
 Figure 7: Ablation study of data filter. We train our model on PIC-98K, PIC-400K and Subject200K respectively.
-Subject-driven Analyses
-Specifically, we discover that different object sizes can lead to inaccurate scores.
-As shown in Figure 6, to avoid the sensitivity, we crop the subject from original image as source images for evaluation.
+4.3 Quantitative Evaluations
+4.3.1 Subject-driven Analyses
+Specifically, we discover that different object sizes can lead to inaccurate scores, which will be detailed in Appendix.
 We compare our proposed PositionIC with several leading methods on DreamBench for both single-subject and multi-subject.
@@
 The evaluation results indicate that PositionIC has remarkable performance on subject consistency and text fidelity.
-Controllable Spacial Generation
+Figure 8: User study for subject-driven generation and data filter.
+(a) User evaluation on Dreambench.
+(b) Filtering consistency of our data filter compared with human in different data quality intervals (from good to bad).
+4.3.2 Controllable Spacial Generation
 Table 3 presents the results of spatial control evaluation.
@@
 For multi-subject evaluation, PositionIC achieves the highest scores in both AP and IoU scores, demonstrating a significant advantage over existing methods.
-User Study
+4.3.3 User Study
 We invite evaluators for an extensive user study.
 For subject-driven generation, we randomly selected 500 images from the results on DreamBench for manual evaluation.
-There are six users scored the results from five dimensions, with each ranging from 0 to 5, and the average score was taken.
+Six users scored the results from five dimensions, each ranging from 0 to 5, and the average score was taken.
 The result presented in Figure 8 (a) shows that our method reaches the highest capability to preserve image features and details while maintaining competitive text adherence.
@@
 As shown in Figure 8 (b), our filter has an average consistency of 0.89 with human annotators.
-Ablation Study
-In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 9, Table 4 and Figure 10.
-Impact of data filter.
+4.4 Ablation Study
+In this section we show the ablation study of our key modules, including data quality and Volumetric Weight Regulation(VWR). More ablation results are presented in the appendix.
+4.4.1 Impact of data filter.
 Results in Figure 7 illustrate the advanced fidelity of BMPDS.
@@
 PIC-400K achieves significantly higher scores than Subject200K, and the filtered data PIC-98K achieves the highest scores overall, which demonstrates the remarkable efficiency of our data synthesis and filtering pipeline.
-[…]
-w/o horizon mask 0.784 0.269 0.686
-w/ horizon mask 0.846 0.269 0.823
-Table 4: Ablation study of horizon mask. Our model perform better subject fidelity on Dreambench after restrict the attention horizon of reference images.
-Figure 10: Qualitative results of ablation study. Horizon mask and our filter pipeline are capable of effective position control and consistent customization.
-Horizon mask.
-As shown in Table 4, CLIP-I and DINO scores significantly drop when training without horizon mask.
-We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistency.
-Conclusion
+Figure 9: Impact of VWR. In the presence of overlaps, VWR is capable of correctly handling the spatial relationships between objects. We set the attention value in overlap region as 1 to disable VWR.
+4.4.2 Abation results of Volumetric Weight Regulation(VWR)
+As shown in Fig.
+9, in Case 1 overlap setting, the doll is instructed to be placed behind the sports car.
+However, the model fails to correctly understand their spatial relationship without VWR.
+In Case 2, VWR is able to prevent the generation of missing objects caused by concept confusion, even when multiple objects overlap.
+5 Conclusion
 In this work, we present PositionIC, an innovative framework capable of customizing multiple subjects with precise position control.
@@
 We hope our work can advance the development of controllable image customization.
-Appendix A Bidirectional Multi-dimensional Perception Data Synthesis
-Detailed Instruction of GPT-4o
+[…]
+Supplementary Material
+Figure 10: Category distribution of PIC-98K.
+In Section 6, we give a detailed description of our data synthesis and filtering pipeline: Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS).
+We further present a selection of example images.
+Section 7 introduces the PositionIC-Bench we use for position evaluation.
+Section  8,9 and 10 provide a detailed discussion of specific experimental setups and additional ablation studies. Furthermore, more example demonstrations can be found in Figure 12.
+6 Bidirectional Multi-dimensional Perception Data Synthesis
+6.1 Detailed Instruction of GPT-4o
 As shown in Figure 15, the message given to GPT-4o consists of Instruction, Evaluation Metric and Response.
@@
 The final score is calculated as the average of all feature scores.
-For descriptions with significant discrepancies, the LLMs is permitted to assign a score of zero to the samples.
+For descriptions with significant discrepancies, the LLMs are permitted to assign a score of zero to the samples.
 Figure 17 demonstrates the samples of Multi-dimensional Perception Data Filter.
@@
 In the second sample, the deer is missing antlers, which resulted in the lowest score on the ”antler” feature.
-Details of PIC-98K Dataset
-We propose PIC-400K utilizing our Bidirectional Multi-dimensional Perception Data Synthesis, a automatic and effective high-consistency data synthesis pipeline.
-Samples of the filtered data PIC-98K is shown in the Figure 13.
+6.2 Details of PIC-98K Dataset
+We propose PIC-400K utilizing our Bidirectional Multi-dimensional Perception Data Synthesis, an automatic and effective high-consistency data synthesis pipeline.
+Samples of the filtered data PIC-98K are shown in the Figure 16.
 BMPDS can synthesize high-fidelity multi-subject images while maintaining high resolution.
@@
 Figure 11: Showcases of PositionIC-Bench.
-Appendix B PositionIC-Bench
+Figure 12: More results about PositionIC.
+7 PositionIC-Bench
 We manually select 252 single-subject samples and 296 multi-subject samples in the benchmark, where the object bounding boxes conform to standard proportions and include challenging positional relationships.
@@
 At the same time, the bounding boxes of smaller objects is appropriately enlarged to obtain more accurate evaluation scores.
+Figure 13: Inaccuracy of directly using CLIP-I and DINO. The revised score is less affected by the size of the subject which can reflect the subject consistency more authentically.
+8 Training Details and Result
+8.1 Setting of Training
+The value of the semantic density $\sigma_{i}$ can be manually set or determined by Eq.
+8.
+In  Eq.
+8, we set the semantic density to a geometric progression with a common ratio of $1+10\lambda$, which represents that $\sigma_{i}$ increases incrementally from far to near to reflect the difference between foreground and background objects.
+$\sigma_{i}=\frac{10\lambda*(1+10\lambda)^{i}}{(1+10\lambda)^{n}-1},$ (8)
+where $n$ is the number of reference images and $i=\{0,1,2,...,n-1\}$ represents the object index ordered from far to near.
+$\lambda$ is a hyperparameter used to control the variation in semantic density.
+A smaller value of $\lambda$ results in less difference in semantic density between objects.
+During training, we set it to 5.
+8.2 More Result
+As shown in Figure 12, we downloaded several foreground images from the internet as additional visual demonstrations, covering four categories: Characters, Items, Animals, and Space.
+In the Characters and Items sections, PositionIC exhibits high fidelity and is capable of generating images with coherent spatial relationships.
+In the Animals section, we demonstrate the plausibility of object interactions under different prompts.
+In the Space section, PositionIC is able to achieve good results even for difficult examples (e.g., generating a planetary ring surrounding the Earth).
+9 Revised Evaluation Metric
+In this section, we elaborate on the calculation methods for our evaluation metrics.
+As shown in Figure 13, directly using the entire image to compute the CLIP or DINO metrics is unreasonable, as both CLIP and DINO compute the similarity of global image features.
+To avoid the sensitivity, we crop the subject from original image as source images for evaluation.
+Figure 14: Qualitative results of ablation study. Visibility-Aware Attention(VAA) and our filter pipeline are capable of effective position control and consistent customization.
+10 More Ablation Study
+In this section, we conduct more detailed ablation studies, including ablations on the data filter and Visibility-Aware Attention(VAA).
+10.1 Impact of Visibility-Aware Attention
+As shown in Figure 4, CLIP-I and DINO scores significantly drop when training without VAA.
+We infer that incorporating a attention mask allows the model to focus more effectively on features in a smaller region rather than globally, which accelerates the convergence and improves the consistency.
+10.2 Qualitative Results
+As shown in Figure 14, we visualize the effect of VAA and data filtering.
+It can be seen that without VAA, the model loses the ability to control position.
+When using unfiltered data, although the data volume is larger, the generated quality is poorer.
+Figure 15: Prompt template of MLLMs in Multi-dimensional Perception Data Filter.
+[…]
+w/o VAA 0.784 0.269 0.686
+w/ VAA 0.846 0.269 0.823
+Table 4: Ablation study of VAA. Our model performs better subject fidelity on DreamBench after restricting the attention area of reference images.
 Figure 16: Showcases of PIC-98K Dataset.
-Figure 14: Prompt template of MLLMs in Multi-dimensional Perception Data Filter.
 Figure 17: Examples of Multi-dimensional Perception Data Filter.

### v6

@@
 Our work represents a significant step towards truly controllable, high-fidelity image customization in multi-entity scenarios.
-Code and data will be publicly released.
+Code and data: https://github.com/MeiGen-AI/PositionIC.
 1 Introduction
-Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images closely aligned with short textual prompts [sdxl, stablediffusion3, ldm, playground, flux].
+Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images closely aligned with short textual prompts [23, 5, 25, 17, 14].
 Beyond generic generation, subject-driven image customization aims to integrate one or more reference objects into user-specified scenes under the guidance of a text prompt.
-Traditionally, this task emphasizes two core goals: (i) preserving the appearance identity of the reference object(s), and (ii) aligning the synthesis with the user’s textual description, where remarkable improvements in visual fidelity have been achieved by recent methods [uno, dreamo, omini_control, wang2025unicombine].
+Traditionally, this task emphasizes two core goals: (i) preserving the appearance identity of the reference object(s), and (ii) aligning the synthesis with the user’s textual description, where remarkable improvements in visual fidelity have been achieved by recent methods [35, 21, 28, 30].
 Figure 1: Results from PositionIC across various controllable image customization tasks.
@@
 Specifically, two coupled bottlenecks remain: (1) the scarcity of large-scale datasets annotated with explicit positional labels of multiple subjects, hindering the learning of spatial reasoning; and (2) the prevalent use of global-level attention mechanisms in diffusion transformers, which inadvertently entangle semantic identity and spatial layout, thereby failing to provide fine-grained, visibility-aware control.
-While a few recent works [ms-diffusion, gligen] have attempted layout control, they typically exhibit a trade-off between spatial precision and subject identity consistency, thus limiting their practical utility.
+While a few recent works [33, 18] have attempted layout control, they typically exhibit a trade-off between spatial precision and subject identity consistency, thus limiting their practical utility.
 To this end, we introduce two tightly coupled components: (i) a scalable Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS) pipeline that generates high-quality, multi-subject, position-annotated data; and (ii) PositionIC, a lightweight layout-aware diffusion framework incorporating a visibility-aware attention mechanism that enables precise, controllable placement of multiple subjects while maintaining identity fidelity.
@@
 A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed.
-Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) [han2024instinctive, bpo], we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks.
+Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) [7, 22], we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks.
 This two-stage filtering markedly improves data reliability.
 Built upon the BMPDS corpus, PositionIC directly integrates our visibility-aware attention mechanism into the diffusion transformer architecture.
-This mechanism is achieved mainly via an operation inspired by NeRF [nerf] termed Volumetric Weight Regulation, which explicitly constructs attention masks based on the likelihood of visibility derived from volumetric rendering principles.
+This mechanism is achieved mainly via an operation inspired by NeRF [20] termed Volumetric Weight Regulation, which explicitly constructs attention masks based on the likelihood of visibility derived from volumetric rendering principles.
 Unlike prevalent global-level attention, our visibility-aware attention mechanism effectively decouples instance-level spatial embeddings from semantic identity features.
@@
 2.1 Subject-driven Generation
-In addition to using text prompts for conditional image generation, current diffusion models [ELITE, li2023blip, huang2024realcustom, ip-adapter, groundingbooth, fastcomposer, MultiDiffusion, feng2025personalize] support reference image input to achieve preservation of subject identity.
-Dreambooth [dreambooth] and LoRa [lora] control the generation of diffusion models through fine-tuning on the specific subject.
-Recently, Diffisuion-Transformers-based subject-driven models [uno, omini_control, omini_control2, dreamo, xiao2025omnigen, flux-kontext] further advance subject-driven generation.
+In addition to using text prompts for conditional image generation, current diffusion models [34, 16, 10, 39, 38, 36, 1, 6] support reference image input to achieve preservation of subject identity.
+Dreambooth [26] and LoRa [9] control the generation of diffusion models through fine-tuning on the specific subject.
+Recently, Diffisuion-Transformers-based subject-driven models [35, 28, 29, 21, 37, 15] further advance subject-driven generation.
 They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation.
 2.2 Position Controllable Generation
-Some works [instancediffusion, regional_flux, wang2024compositional, jimenez2023mixture, MultiDiffusion, shi2025layoutcot, LaRender] have attempted to generate with precise layout control.
-Gligen [gligen] encodes Fourier embedding as grounding tokens to inject position information.
-MS-diffusion [ms-diffusion] utilizes a grounding resampler correlating visual information with specific entities and spatial constraints.
+Some works [32, 4, 31, 12, 1, 27, 40] have attempted to generate with precise layout control.
+Gligen [18] encodes Fourier embedding as grounding tokens to inject position information.
+MS-diffusion [33] utilizes a grounding resampler correlating visual information with specific entities and spatial constraints.
 However, they still suffer from inconsistencies in vision and position, which hinder further application.
@@
 Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals.
-However, existing open-source datasets such as Subject200K [omini_control] generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information.
+However, existing open-source datasets such as Subject200K [28] generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information.
 We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations.
@@
 3.1.2 Multi-dimensional Perception Data Filter
-Previous studies [cao2024mmfuser, he2025analyzing] have shown that MLLMs have limited capability in recognizing fine-grained details in images.
+Previous studies [2, 8] have shown that MLLMs have limited capability in recognizing fine-grained details in images.
 Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering.
 Specifically, we divide the filtering process into three levels based on granularity.
-Firstly, we segment the subjects from data pairs and utilize CLIP-I [CLIP] and DINO [DINO] scores $s_{v}$ to filter out images with significantly lower consistency.
+Firstly, we segment the subjects from data pairs and utilize CLIP-I [24] and DINO [3] scores $s_{v}$ to filter out images with significantly lower consistency.
 Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color and so on.
-Lastly, we employ Describe Anything Model (DAM) [describe_anything], a description expert to obtain detailed textual descriptions for each subject.
+Lastly, we employ Describe Anything Model (DAM) [19], a description expert to obtain detailed textual descriptions for each subject.
 Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@
 As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text cross attention.
-If extending to single or multi-subject custom generation via concatenating method like [uno, omini_control, flux-kontext, eligen], the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
+If extending to single or multi-subject custom generation via concatenating method like [35, 28, 15, 41], the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
 To achieve effective attention accumulation, we propose Visibility-Aware Attention, which utilizes volume rendering(VR) to explicitly define the area that can be focused on for each reference image.
-Previous works [eligen, regional_flux, zhang2025creatidesign, LaRender] explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
+Previous works [41, 4, 42, 40] explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
 The overall mechanism is shown in Figure 3.
 3.2.1 Volume Rendering(VR)
-The weighted mask utilizes VR [volume_render] to handle the occlusion relationships between individual objects.
-Similar to NeRF [nerf], the expected color $C(r)$ observed by the camera along the distance from the far bound $t_{f}$ to the near bound $t_{n}$ is:
+The weighted mask utilizes VR [13] to handle the occlusion relationships between individual objects.
+Similar to NeRF [20], the expected color $C(r)$ observed by the camera along the distance from the far bound $t_{f}$ to the near bound $t_{n}$ is:
 $C(r)=\int_{t_{n}}^{t_{f}}T(t)\sigma(r(t))c(r(t),d)dt,$
@@
 4.1.2 Evaluation Metrics
-For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench [dreambooth].
+For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench [26].
 For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
 For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench.
-We use Vision-R1 [visionr1] to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
+We use Vision-R1 [11] to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
 Method Single-Subject Multi-Subject
 […]
-RPF [regional_flux] 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
-MS-Diffusion [ms-diffusion] 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
-Instance-Diffusion [instancediffusion] 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
-Gligen [gligen] 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
+RPF [4] 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
+MS-Diffusion [33] 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
+Instance-Diffusion [32] 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
+Gligen [18] 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
 PositionIC(Ours)({Ours}) 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
@@
 We hope our work can advance the development of controllable image customization.
-[…]
 Supplementary Material